An airline estimates that 95% of the people booked on their flights actually show up. If the airline books 62 people on a flight for which the maximum number is 60, what is the probability that the number of people who show up will exceed the capacity of the plane?
Answer by stanbon(48568)


An airline estimates that 95% of the people booked on their flights actually show up. If the airline books 62 people on a flight for which the maximum number is 60, what is the probability that the number of people who show up will exceed the capacity of the plane?
---------------
P(X=61)+P(X=62)
=62C61(0.95)^61*(.05)^1 + 62C62(0.95)^62*(0.05)^0
=0.136 + 0,0416
=0.177
Cheers,
Stan H.